one person is to be selected at random from of group of 25 people.the probility that seleact person will ne a male is 0.44 and the probility that selected person will be a male who was born after the 1960 is 0.28 r


1
Expert's answer
2021-05-07T10:38:47-0400

We have total number of people = 25.


Now, number of mean in the group "= 25 \times 0.44 = 11"


Since the probability of selecting men "= 0.44 = \dfrac{11}{25}"


Similarly number of men born 1960 "= 25 \times 0.28 = 7"


The number of males in the group who were born in 1960 or later "= 11-7 = 4"